OpenStudy (anonymous):

What is the solution to x + 11.4 = 13.9?

jhonyy9 (jhonyy9):

2,5

OpenStudy (liliakarina):

2.5

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

x + 11.4 = 13.9 x = 13.9 - 11.4 x = 2.5

OpenStudy (anonymous):

x=2.5